Calculate the number of molecules of sulphur (S8) present in 16 g of solid sulphur.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Molar mass of S8 sulphur = 256 g = 6.022 × 1023 molecule
Given mass of sulphur = 16 g
Molecules of sulphur =16 x 6.022 x 10^23
256
= 96.35 x 10^23
256

= 0.376 × 10^23
= 3.76 × 10^22molecules

Thanks